Use the normal distribution of SAT critical reading scores for which the mean is 510 and the standard deviation is 118. Assume the variable x is normally distributed (a) What percent of the SAT verbal scores are less than 600? (b) If 1000 SAT verbal scores are randomly selected, about how many would you expect to be greater than 575? (a) Approximately%of the SAT verbal scores are less than 600. (Round to two decimal places as needed.) (b) You would expect that approximately SAT verbal scores would be greater than 575. (Round to the nearest whole number as needed.)
Answer #1

Given,

\mu = 510 ,  \sigma = 118

We convert this to standard normal as

P( X < x) = P( Z < x - \mu / \sigma )

a)

P( X < 600) = P(Z < 600 - 510 / 118 )

= P( Z < 0.7627)

= 0.7772

b)

P( X > 575) = P( Z > 575 - 510 / 118 )

= P( Z > 0.5508)

= 1 - P( Z < 0.5508)

= 1 - 0.7091

= 0.2909

So among the 1000 scores, we expect 0.2909 * 1000 = 291